OpenStudy (anonymous):

Graph the line with slope 3 passing through the point (1,-4)

OpenStudy (anonymous):

Use the formula \[(y-y1) = m(x-x1)\] where m = slopes

OpenStudy (anonymous):

Once you get the equation you can make a table and find coordinates to graph

OpenStudy (anonymous):

Or you could plot the point (1,-4) then use the slope to find several other points and draw a line through them. Since the slope is 3 that would mean you go up 3 and right 1 OR down 3 and left 1 to get to a new point.

OpenStudy (mertsj):

put a dot on (1,-4). Put your pencil on that point and then follow the directions given by the slope 3 which is 3 over 1 which means go up three and then go right 1. Make a dot. Two points determine a line. Draw the line.